# Self-testing of genuine multipartite non-local and non-maximally entangled states

### v2

@@ Abstract
Self-testing enables the characterization of quantum systems with minimal assumptions their internal working. As such it represents the strongest form of certification for quantum systems. In the existing self-testing literature, self-testing states which are not maximally entangled, but exhibit genuine multipartite nonlocality, have remained an open problem. This is particularly important because, for many-body systems, genuine multipartite nonlocality has been recognized as the strongest form of multipartite quantum correlation. In this work, we present a Hardy-like paradox for scenarios involving arbitrary number of parties. This paradox is a tool for detecting genuine multipartite nonlocality, allowing for the specific identification and self-testing of states that defy the paradox’s limits the most, which turn out to be non-maximally multipartite entangled states. While recent results [Šupić et al., Nature Physics, 2023] suggest network self-testing as a means to self-test all quantum states, here we operate within the standard self-testing framework to self-test genuine multipartite non-local and non-maximally entangled states.
@@
In this paper, we answer the above question in the affirmative. We present a Hardy-like paradox for a scenario involving an arbitrary number of participants, each having access to two measurements that can yield any arbitrary number of outcomes. Our paradox helps witness genuine multipartite nonlocality. The state maximally violating the constraints of our paradox can be self-tested. Regarding robust self-testing, we provide a partial robustness result in the three-party scenario where we consider the non-ideal state to be a three-qubit state. Moreover, it is important to note that the aforementioned state is non-maximally entangled. We highlight that unlike Ref [32], our work does not require quantum networks. We operate in the standard Bell-test based self-testing framework and do not need any additional assumptions unlike Ref [32].
@@
The discrete random variable representing the common cause is denoted as $\lambda$, with a distribution $\nu(\lambda)\geq 0$ and the normalization condition $\sum_{\lambda}\nu(\lambda)=1$. The conditional probability distribution $P_{A_{i}}(x_{i}|\hat{x}_{i},\lambda)$ characterizes the behavior of party $\mathcal{A}_{i}$. A distribution $P(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})$ is considered non-local if it cannot be decomposed as described in equation 1. Given the measurement setups, we work under the assumption of the NS principle [37]. This implies that party $\mathcal{A}_{1}$ is unable to communicate with other parties through the choice of measurement. Specifically, the equality $P_{A_{2}A_{3}}(x_{2}x_{3}|\hat{x}_{2}\hat{x}_{3})=P_{A_{2}A_{3}}(x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})=\sum_{x_{1}}P(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3}),\forall x_{1}$ holds, and the analogous statements apply to parties $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$.
@@
with $\nu_{i}(\lambda_{i})\geq 0$ and $\sum_{i,\lambda_{i}}\nu_{i}(\lambda_{i})=1$. Distributions denoted as $P(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})$ that cannot be expressed in the form described in equation 2 are termed as genuine tripartite non-local distributions. According to findings in [38], the original concept of multipartite non-locality introduced by Svetlichny [34] encounters practical challenges when operationalized. To address these challenges, it is assumed that the NS principle [39] also holds at the level of distributions $P_{A_{i}A_{j}}(x_{i}x_{j}|\hat{x}_{i}\hat{x}_{j},\lambda)$, implying that the conditional probabilities $P(x_{i}|\hat{x}_{i},\lambda)=P(x_{i}|\hat{x}_{i}\hat{x}_{j},\lambda)=\sum_{x_{j}}P(x_{i}x_{j}|\hat{x}_{i}\hat{x}_{j},\lambda),\forall\hat{x}_{j}$, are well defined for all $\lambda$.
@@
If a non-local correlation satisfies the conditions in 3 with $\mathcal{S}_{3}>0$, then it is a genuine tripartite non-local correlation.
@@
We will first show that no local theory can satisfy conditions 3 with $\mathcal{S}_{3}>0$. Let us define a local hidden variable $\lambda$ that captures the full description of the system. This $\lambda$ takes values from a set $\Omega$. Moreover, we assert that the comprehensive depiction of the joint system is denoted by $\nu(\lambda)$. In a local description, the joint probability can be expressed as follows: $P(x_{1},x_{2},x_{3}|\hat{x}_{1},\hat{x}_{2},\hat{x}_{3})=\sum_{\lambda\in\Omega}\nu(\lambda)\prod_{j=1}^{3}p(x_{j}|\hat{x}_{j},\lambda)$, where $\nu(\lambda)\geq 0$, $\sum_{\lambda}\nu(\lambda)=1$. Let’s assume that $\mathcal{S}_{3}>0$, which indicates that $p_{3}$ is greater than 00. This implies the existence of a nonempty subset $\Lambda\subseteq\Omega$, such that for all $\lambda\in\Lambda$, $P(+|A_{0},\lambda)=P(+|B_{0},\lambda)=P(+|C_{0},\lambda)=1$. By equation 3 we have $P(-|A_{1},\lambda)=P(-|B_{1},\lambda)=P(-|C_{1},\lambda)=1$. Therefore, we can deduce that $P(-,-,-|A_{1},B_{1},C_{1},\Lambda)=P(+,+,+|A_{0},B_{0},C_{0})$. However, since $\Lambda\subseteq\Omega$, it follows that $P(-,-,-|A_{1},B_{1},C_{1},\Lambda)\leq P(-,-,-|A_{1},B_{1},C_{1})$. In other words, $\mathcal{S}_{3}\leq 0$, which contradicts the initial assumption of $\mathcal{S}_{3}>0$. Therefore, we can conclude that if any correlation satisfies 3 and gives $\mathcal{S}_{3}>0$, then it must be non-local.
Now we will show that the non-locality we have is a genuine one. If this correlation were not genuine, it could be broken down according to the expression provided in equation 2. Given that $\mathcal{S}_{3}>0$, at least one component within the decomposition must yield a non-zero value for $P(+++|A_{0}B_{0}C_{0})$. Let’s consider this specific component to be the initial one: $P_{AB}(x_{1}x_{2}|\hat{x}_{1}\hat{x}_{2},\lambda_{1})P_{C}(x_{3}|\hat{x}_{3},\lambda_{1})$. Consequently, $P_{AB}(++|A_{0}B_{0},\lambda_{1})P_{C}(+|C_{0},\lambda_{1})\neq 0$. The condition $P_{BC}(++|B_{1}C_{0})=0$ leads to the conclusion that $P_{B}(+|B_{1},\lambda_{1})P_{C}(+|C_{0},\lambda_{1})$ must be zero. This, in turn, implies that $P_{B}(+|B_{1},\lambda_{1})=0$, given that $P_{C}(+|C_{0},\lambda_{1})$ cannot be zero. Therefore, we deduce that $P_{B}(-|B_{1},\lambda_{1})=1$. Consequently, party $B$ can jointly measure both $B_{0}$ and $B_{1}$. By utilizing the findings from [40] that joint measureability within any no-signaling theory never violates a Bell inequality, we can ascertain that $P_{AB}(x_{1}x_{2}|\hat{x}_{1}\hat{x}_{2},\lambda_{1})=P_{A}(x_{1}|\hat{x}_{1},\lambda_{1})P_{B}(x_{2}|\hat{x}_{2},\lambda_{1})$. Similarly, all the other components within the decomposition become fully product-based. As a result, the distributions $P_{A_{1}A_{2}A_{3}}(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})$ transform into fully local distribution, which contradicts the initial premise considering $\mathcal{S}_{3}>0$. ∎
@@
We consider the subsystem $\mathcal{A}_{N+1}$ to be $\mathcal{A}_{1}$. Similar to the tripartite case, we define $p_{N}\equiv P(1,1,\dots,1|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})$, $q_{N}\equiv P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N})$ and their difference to be $\mathcal{S}_{N}:=p_{N}-q_{N}$. We will show that only non-local correlation can satisfy 4 with $\mathcal{S}_{N}>0$ and furthermore, this non-locality is genuine. The following theorem is motivated from Ref [41].
@@
If a non-local correlation satisfies the conditions in 4 with $\mathcal{S}_{N}>0$, then it is a genuine multipartite non-local correlation.
@@
We will first show that no local theory can satisfy conditions 4 with $\mathcal{S}_{N}>0$. Let us define a local hidden variable $\lambda$ that captures the full description of the system. This $\lambda$ takes values from a set $\Omega$. Moreover, we assert that the comprehensive depiction of the joint system is denoted by $\nu(\lambda)$. In a local description, there exist conditional probabilities $p_{A_{j}}(u_{j}|\hat{u}_{j},\lambda)$ and $p_{A_{j}}(v_{j}|\hat{v}_{j},\lambda)$ such that the joint probability can be expressed as follows: $P(x_{1},x_{2},\dots,x_{N}|\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N})=\sum_{\lambda\in\Omega}\nu(\lambda)\prod_{j=1}^{N}p_{A_{j}}(x_{j}|\hat{x}_{j},\lambda)$, where $\hat{x}_{j}\in\{\hat{u}_{j},\hat{v}_{j}\}$. Now, let’s assume that $\mathcal{S}_{N}>0$, indicating that $p_{N}>0$. This implies the existence of a nonempty subset $\Lambda\subseteq\Omega$, such that for all $\lambda\in\Lambda$, $P_{A_{i}}(1|\hat{u}_{i},\lambda)=1$ for all $i$. Consequently, for all $r$ where $v_{r}\neq d_{i}$, we have $P_{A_{i}}(v_{r}|\hat{v}_{i},\lambda)=0$, which further implies $P(d_{i}|\hat{v}_{i},\lambda)=1$. Therefore, we can deduce that $P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N},\Lambda)=P(1,1,\dots,1|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})=p_{N}$. However, since $\Lambda\subseteq\Omega$, it follows that $P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N},\Lambda)\leq P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N})$. In other words, $\mathcal{S}_{N}\leq 0$, which contradicts the initial assumption of $\mathcal{S}_{N}$. Therefore, we can conclude that if any correlation satisfies 4 and gives $\mathcal{S}_{N}>0$, then it must be non-local.
@@
Now we will show the genuineness of the non-locality. Consider a $N$-partite probability distribution $P(x_{1},x_{2},\dots,x_{N}|\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N})$ satisfying conditions in 4 with $\mathcal{S}_{N}>0$, which is not genuine. Hence, it is a convex mixture of bi-local correlation. Let’s consider a partition $(1,2,\dots,m)$ versus $(m+1,m+2,\dots,N)$. Consequently, there must be at least one element within the convex combination that contributes non-zero value to the condition $P(1,1,\dots,1|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})>0$, given $p_{N}>q_{N}\geq 0$. Let’s assume such an element takes the form $P_{1}(x_{1},\dots,x_{m}|\hat{x}_{1},\dots,\hat{x}_{m})P_{2}(x_{m+1},\dots,x_{N}|\hat{x}_{m+1},\dots,\hat{x}_{N})$. As $p_{N}>0$ we must have,
$\begin{split}P_{1}(1,\dots,1|\hat{u}_{1},\dots,\hat{u}_{m})P_{2}(1,\dots,1|\hat{u}_{m+1},\dots,\hat{u}_{N})>0\end{split}$ (5)
whereas the equation 4 imply two important relations,
$\begin{split}&P_{1}(v_{m}\neq d_{m}|\hat{v}_{m})P_{2}(1|\hat{u}_{m+1})=0,\\ &P_{2}(v_{N}\neq d_{N}|\hat{v}_{N})P_{1}(1|\hat{u}_{1})=0.\end{split}$ (6)
@@
The form of $P_{1}$: Since by 5, one definitely have $P_{2}(1|\hat{u}_{m+1})\neq 0$, we get $P_{1}(v_{m}\neq d_{m}|\hat{v}_{m})=0$, hence $P_{1}(d_{m}|\hat{v}_{m})=1$. By utilizing the findings from [40] that joint measureability within any no-signaling theory never violates a Bell inequality, $P_{1}(x_{1},\dots,x_{m}|\hat{x}_{1},\dots,\hat{x}_{m})$ factorizes to certain $P^{\prime}_{1}(x_{1},\dots,x_{m-1}|\hat{x}_{1},\dots,\hat{x}_{m-1})P^{m}_{1}(x_{m}|\hat{x}_{m})$. The conditions in 4 also implies that $P^{\prime}_{1}(v_{m-1}\neq d_{m-1}|\hat{v}_{m-1})P^{m}_{1}(1|\hat{u}_{m})=0$, however again since $P^{m}_{1}(1|\hat{u}_{m})\neq 0$ one has $P^{\prime}_{1}(d_{m-1}|\hat{v}_{m-1})=1$. Using the same argument it is easy to see that $P_{1}$ fully factorizes.
@@
The form of $P_{2}$: We have $P_{2}(v_{N}\neq d_{N}|\hat{v}_{N})P_{1}(1|\hat{u}_{1})=0$, from which $P_{2}(v_{N}\neq d_{N}|\hat{v}_{N})=0$, thus $P_{2}(d_{N}|\hat{v}_{N})=1$. Hence $P_{2}$ also fully factories as $P_{1}$.
@@
Thus the distribution $P(x_{1},x_{2},\dots,x_{N}|\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N})$ become fully local. Such probability distribution cannot satisfy the equation 4 for $\mathcal{S}_{N}>0$. Hence we are done. Since the proof is the same for all cuts, no mixture of bi-local probability distribution for different cuts can satisfy all the conditions. ∎
@@
First, we demonstrate self-testing for the tripartite case, then extend to the $N$-partite scenario. Alice, Bob, and Charlie each choose measurements $\hat{x}_{1}\in\{A_{0},A_{1}\}$, $\hat{x}_{2}\in\{B_{0},B_{1}\}$, and $\hat{x}_{3}\in\{C_{0},C_{1}\}$, respectively. Their binary outcomes are denoted as $x_{1}\in\{\pm 1\}$, $x_{2}\in\{\pm 1\}$, and $x_{3}\in\{\pm 1\}$. To illustrate genuine non-locality, consider that Alice, Bob, and Charlie share the following general pure three-qubit state,
@@
The class of pure three-qubit state that satisfies the condition 3 is of the form:
@@
$P(x,y,z|\hat{x},\hat{y},\hat{z})=\operatorname{Tr}(\rho\Pi_{x|\hat{x}}\otimes\Pi_{y|\hat{y}}\otimes\Pi_{z|\hat{z}}).$ (10)
@@
$\hat{x}$ $=(+1)\Pi_{+|\hat{x}}+(-1)\Pi_{-|\hat{x}},\quad\hat{x}\in\{A_{0},A_{1}\}$ (11)
$\hat{y}$ $=(+1)\Pi_{+|\hat{y}}+(-1)\Pi_{-|\hat{y}},\quad\hat{y}\in\{B_{0},B_{1}\}$
$\hat{z}$ $=(+1)\Pi_{+|\hat{z}}+(-1)\Pi_{-|\hat{z}},\quad\hat{z}\in\{C_{0},C_{1}\}$

### v3

@@ Abstract
Self-testing enables the characterization of quantum systems with minimal assumptions on their internal working as such it represents the strongest form of certification for quantum systems. In the existing self-testing literature, self-testing states that are not maximally entangled, but exhibit genuine multipartite nonlocality, have remained an open problem. This is particularly important because, for many-body systems, genuine multipartite nonlocality has been recognized as the strongest form of multipartite quantum correlation. In this work, we present a Cabello-like paradox for scenarios involving an arbitrary number of parties. This paradox is a tool for detecting genuine multipartite nonlocality, allowing for the specific identification and self-testing of states that defy the paradox’s limits the most, which turn out to be non-maximally multipartite entangled states. While recent results [Šupić et al., Nature Physics, 2023] suggest network self-testing as a means to self-test all quantum states, here we operate within the standard self-testing framework to self-test genuine multipartite non-local and non-maximally entangled states.
@@
In this paper, we answer the above question in the affirmative. We present a Cabello-like paradox for a scenario involving an arbitrary number of participants, each having access to two measurements that can yield any arbitrary number of outcomes. Our paradox helps witness genuine multipartite nonlocality. The state maximally violating the constraints of our paradox can be self-tested. Regarding robust self-testing, we provide a partial robustness result in the three-party scenario where we consider the non-ideal state to be a three-qubit state. Moreover, it is important to note that the aforementioned state is non-maximally entangled. We highlight that unlike Ref [32], our work does not require quantum networks. We operate in the standard Bell-test based self-testing framework and do not need any additional assumptions unlike Ref [32].
@@
The discrete random variable representing the common cause is denoted as $\lambda$, with a distribution $\nu(\lambda)\geq 0$ and the normalization condition $\sum_{\lambda}\nu(\lambda)=1$. The conditional probability distribution $P_{A_{i}}(x_{i}|\hat{x}_{i},\lambda)$ characterizes the behavior of party $\mathcal{A}_{i}$. A distribution $P(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})$ is considered non-local if it cannot be decomposed as described in equation (1). Given the measurement setups, we work under the assumption of the NS principle [37]. This implies that party $\mathcal{A}_{1}$ is unable to communicate with other parties through the choice of measurement. Specifically, the equality $P_{A_{2}A_{3}}(x_{2}x_{3}|\hat{x}_{2}\hat{x}_{3})=P_{A_{2}A_{3}}(x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})=\sum_{x_{1}}P(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3}),\forall x_{1}$ holds, and the analogous statements apply to parties $\mathcal{A}_{2}$ and $\mathcal{A}_{3}$.
@@
with $\nu_{i}(\lambda_{i})\geq 0$ and $\sum_{i,\lambda_{i}}\nu_{i}(\lambda_{i})=1$. Distributions denoted as $P(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})$ that cannot be expressed in the form described in equation (2) are termed as genuine tripartite non-local distributions. According to findings in [38], the original concept of multipartite non-locality introduced by Svetlichny [34] encounters practical challenges when operationalized. To address these challenges, it is assumed that the NS principle [39] also holds at the level of distributions $P_{A_{i}A_{j}}(x_{i}x_{j}|\hat{x}_{i}\hat{x}_{j},\lambda)$, implying that the conditional probabilities $P(x_{i}|\hat{x}_{i},\lambda)=P(x_{i}|\hat{x}_{i}\hat{x}_{j},\lambda)=\sum_{x_{j}}P(x_{i}x_{j}|\hat{x}_{i}\hat{x}_{j},\lambda),\forall\hat{x}_{j}$, are well defined for all $\lambda$.
@@
If a non-local correlation satisfies the conditions in (3) with $\mathcal{S}_{3}>0$, then it is a genuine tripartite non-local correlation.
@@
We will first show that no local theory can satisfy conditions (3) with $\mathcal{S}_{3}>0$. Let us define a local hidden variable $\lambda$ that captures the full description of the system. This $\lambda$ takes values from a set $\Omega$. Moreover, we assert that the comprehensive depiction of the joint system is denoted by $\nu(\lambda)$. In a local description, the joint probability can be expressed as follows: $P(x_{1},x_{2},x_{3}|\hat{x}_{1},\hat{x}_{2},\hat{x}_{3})=\sum_{\lambda\in\Omega}\nu(\lambda)\prod_{j=1}^{3}p(x_{j}|\hat{x}_{j},\lambda)$, where $\nu(\lambda)\geq 0$, $\sum_{\lambda}\nu(\lambda)=1$. Let’s assume that $\mathcal{S}_{3}>0$, which indicates that $p_{3}$ is greater than 00. This implies the existence of a nonempty subset $\Lambda\subseteq\Omega$, such that for all $\lambda\in\Lambda$, $P(+|A_{0},\lambda)=P(+|B_{0},\lambda)=P(+|C_{0},\lambda)=1$. By equation (3) we have $P(-|A_{1},\lambda)=P(-|B_{1},\lambda)=P(-|C_{1},\lambda)=1$. Therefore, we can deduce that $P(-,-,-|A_{1},B_{1},C_{1},\Lambda)=P(+,+,+|A_{0},B_{0},C_{0})$. However, since $\Lambda\subseteq\Omega$, it follows that $P(-,-,-|A_{1},B_{1},C_{1},\Lambda)\leq P(-,-,-|A_{1},B_{1},C_{1})$. In other words, $\mathcal{S}_{3}\leq 0$, which contradicts the initial assumption of $\mathcal{S}_{3}>0$. Therefore, we can conclude that if any correlation satisfies (3) and gives $\mathcal{S}_{3}>0$, then it must be non-local.
Now we will show that the non-locality we have is a genuine one. If this correlation were not genuine, it could be broken down according to the expression provided in equation (2). Given that $\mathcal{S}_{3}>0$, at least one component within the decomposition must yield a non-zero value for $P(+++|A_{0}B_{0}C_{0})$. Let’s consider this specific component to be the initial one: $P_{AB}(x_{1}x_{2}|\hat{x}_{1}\hat{x}_{2},\lambda_{1})P_{C}(x_{3}|\hat{x}_{3},\lambda_{1})$. Consequently, $P_{AB}(++|A_{0}B_{0},\lambda_{1})P_{C}(+|C_{0},\lambda_{1})\neq 0$. The condition $P_{BC}(++|B_{1}C_{0})=0$ in equation (3) leads to the conclusion that $P_{B}(+|B_{1},\lambda_{1})P_{C}(+|C_{0},\lambda_{1})$ must be zero. This, in turn, implies that $P_{B}(+|B_{1},\lambda_{1})=0$, given that $P_{C}(+|C_{0},\lambda_{1})$ cannot be zero. Therefore, we deduce that $P_{B}(-|B_{1},\lambda_{1})=1$. Now as the NS principle is assumed at the level of distribution $P_{AB}(x_{1}x_{2}|\hat{x}_{1}\hat{x}_{2},\lambda_{1})$, a deterministic marginal of $P_{AB}(x_{1}x_{2}|\hat{x}_{1}\hat{x}_{2},\lambda_{1})$ indicates that $P_{AB}(x_{1}x_{2}|\hat{x}_{1}\hat{x}_{2},\lambda_{1})=P_{A}(x_{1}|\hat{x}_{1},\lambda_{1})P_{B}(x_{2}|\hat{x}_{2},\lambda_{1})$. Similarly, all the other components within the decomposition become fully product-based. As a result, the distributions $P_{ABC}(x_{1}x_{2}x_{3}|\hat{x}_{1}\hat{x}_{2}\hat{x}_{3})$ transform into fully local distribution, which contradicts the initial premise considering $\mathcal{S}_{3}>0$. ∎
@@
We consider the subsystem $\mathcal{A}_{N+1}$ to be $\mathcal{A}_{1}$. Similar to the tripartite case, we define $p_{N}\equiv P(1,1,\dots,1|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})$, $q_{N}\equiv P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N})$ and their difference to be $\mathcal{S}_{N}:=p_{N}-q_{N}$. We will show that only non-local correlation can satisfy (4) with $\mathcal{S}_{N}>0$ and furthermore, this non-locality is genuine.
@@
If a non-local correlation satisfies the conditions in (4) with $\mathcal{S}_{N}>0$, then it is a genuine multipartite non-local correlation.
@@
We will first show that no local theory can satisfy conditions (4) with $\mathcal{S}_{N}>0$. Let us define a local hidden variable $\lambda$ that captures the full description of the system. This $\lambda$ takes values from a set $\Omega$. Moreover, we assert that the comprehensive depiction of the joint system is denoted by $\nu(\lambda)$. In a local description, there exist conditional probabilities $p_{A_{j}}(u_{j}|\hat{u}_{j},\lambda)$ and $p_{A_{j}}(v_{j}|\hat{v}_{j},\lambda)$ such that the joint probability can be expressed as follows: $P(x_{1},x_{2},\dots,x_{N}|\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N})=\sum_{\lambda\in\Omega}\nu(\lambda)\prod_{j=1}^{N}p_{A_{j}}(x_{j}|\hat{x}_{j},\lambda)$, where $\hat{x}_{j}\in\{\hat{u}_{j},\hat{v}_{j}\}$. Now, let’s assume that $\mathcal{S}_{N}>0$, indicating that $p_{N}>0$. This implies the existence of a nonempty subset $\Lambda\subseteq\Omega$, such that for all $\lambda\in\Lambda$, $P_{A_{i}}(1|\hat{u}_{i},\lambda)=1$ for all $i$. Consequently, for all $r$ where $v_{r}\neq d_{i}$, we have $P_{A_{i}}(v_{r}|\hat{v}_{i},\lambda)=0$, which further implies $P(d_{i}|\hat{v}_{i},\lambda)=1$. Therefore, we can deduce that $P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N},\Lambda)=P(1,1,\dots,1|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})=p_{N}$. However, since $\Lambda\subseteq\Omega$, it follows that $P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N},\Lambda)\leq P(d_{1},d_{2},\dots,d_{N}|\hat{v}_{1},\hat{v}_{2},\dots,\hat{v}_{N})$. In other words, $\mathcal{S}_{N}\leq 0$, which contradicts the initial assumption of $\mathcal{S}_{N}$. Therefore, we can conclude that if any correlation satisfies (4) and gives $\mathcal{S}_{N}>0$, then it must be non-local.
@@
Now we will show the genuineness of the non-locality. Consider a $N$-partite probability distribution $P(x_{1},x_{2},\dots,x_{N}|\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N})$ satisfying conditions in (4) with $\mathcal{S}_{N}>0$, which is not genuine. Hence, it is a convex mixture of bi-local correlation. Let’s consider a partition $(1,2,\dots,m)$ versus $(m+1,m+2,\dots,N)$. Consequently, there must be at least one element within the convex combination that contributes non-zero value to the condition $P(1,1,\dots,1|\hat{u}_{1},\hat{u}_{2},\dots,\hat{u}_{N})>0$, given $p_{N}>q_{N}\geq 0$. Let’s assume such an element takes the form $P_{1}(x_{1},\dots,x_{m}|\hat{x}_{1},\dots,\hat{x}_{m},\lambda)P_{2}(x_{m+1},\dots,x_{N}|\hat{x}_{m+1},\dots,\hat{x}_{N},\lambda)$. As $p_{N}>0$ we must have,
$\begin{split}P_{1}(1,\dots,1|\hat{u}_{1},\dots,\hat{u}_{m},\lambda)P_{2}(1,\dots,1|\hat{u}_{m+1},\dots,\hat{u}_{N},\lambda)>0\end{split}$ (5)
From the equation (4), we have $P_{A_{m}A_{m+1}}(v_{m}\neq d_{m},1|\hat{v}_{m},\hat{u}_{m+1})=0$ and $P_{A_{N}A_{1}}(v_{N}\neq d_{N},1|\hat{v}_{N},\hat{u}_{1})=0$. Hence, we have two important relations,
$\begin{split}&P_{1}(v_{m}\neq d_{m}|\hat{v}_{m},\lambda)P_{2}(1|\hat{u}_{m+1},\lambda)=0,\\ &P_{2}(v_{N}\neq d_{N}|\hat{v}_{N},\lambda)P_{1}(1|\hat{u}_{1},\lambda)=0.\end{split}$ (6)
The form of $P_{1}$: Since by (5), we have $P_{2}(1,\dots,1|\hat{u}_{m+1},\dots,\hat{u}_{N},\lambda)\neq 0$, one definitely have $P_{2}(1|\hat{u}_{m+1},\lambda)\neq 0$. We also get $P_{1}(v_{m}\neq d_{m}|\hat{v}_{m},\lambda)P_{2}(1|\hat{u}_{m+1},\lambda)=0$ from equation (6). Hence, we get $P_{1}(v_{m}\neq d_{m}|\hat{v}_{m},\lambda)=0$ and $P_{1}(d_{m}|\hat{v}_{m},\lambda)=1$. Now as the NS principle is assumed at the level of distribution $P_{1}(x_{1}x_{2}\dots x_{m}|\hat{x}_{1}\hat{x}_{2}\dots\hat{x}_{m},\lambda)$, a deterministic marginal of $P_{1}(x_{1}x_{2}\dots x_{m}|\hat{x}_{1}\hat{x}_{2}\dots\hat{x}_{m},\lambda)$ indicates that $P_{1}(x_{1}x_{2}\dots x_{m}|\hat{x}_{1}\hat{x}_{2}\dots\hat{x}_{m},\lambda)$ factorizes to certain $P^{\prime}_{1}(x_{1},\dots,x_{m-1}|\hat{x}_{1},\dots,\hat{x}_{m-1},\lambda)P^{m}_{1}(x_{m}|\hat{x}_{m},\lambda)$.
@@
From the equation (4), we also have $P_{A_{m-1}A_{m}}(v_{m-1}\neq d_{m-1},1|\hat{v}_{m-1},\hat{u}_{m})=0$. As $P_{1}(x_{1}x_{2}\dots x_{m}|\hat{x}_{1}\hat{x}_{2}\dots\hat{x}_{m},\lambda)$ factorizes we have $P^{\prime}_{1}(v_{m-1}\neq d_{m-1}|\hat{v}_{m-1},\lambda)P^{m}_{1}(1|\hat{u}_{m},\lambda)=0$, however again using equation (5) we have $P^{m}_{1}(1|\hat{u}_{m},\lambda)\neq 0$ and it implies that $P^{\prime}_{1}(d_{m-1}|\hat{v}_{m-1},\lambda)=1$. Using the same argument as above $P^{\prime}_{1}$ factorizes. Hence eventually it is easy to see that $P_{1}$ fully factorizes.
@@
The form of $P_{2}$: We have $P_{1}(1|\hat{u}_{1},\lambda)\neq 0$ and $P_{2}(v_{N}\neq d_{N}|\hat{v}_{N},\lambda)P_{1}(1|\hat{u}_{1},\lambda)=0$ from equation (5) and (6) respectively. Hence we have $P_{2}(v_{N}\neq d_{N}|\hat{v}_{N},\lambda)=0$, thus $P_{2}(d_{N}|\hat{v}_{N},\lambda)=1$. Using the same argument for factorization of $P_{1}$, $P_{2}$ also fully factorizes.
@@
Thus the distribution $P_{1}(x_{1},\dots,x_{m}|\hat{x}_{1},\dots,\hat{x}_{m},\lambda)P_{2}(x_{m+1},\dots,x_{N}|\hat{x}_{m+1},\dots,\hat{x}_{N},\lambda)$ fully factorizes. Since the proof is the same for all cuts, the distribution $P(x_{1},x_{2},\dots,x_{N}|\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N})$ becomes fully local. Such probability distribution cannot satisfy the equation (4) for $\mathcal{S}_{N}>0$. Hence we are done. ∎
@@
We will now demonstrate self-testing for the tripartite case. Alice, Bob, and Charlie each choose measurements $\hat{x}_{1}\in\{A_{0},A_{1}\}$, $\hat{x}_{2}\in\{B_{0},B_{1}\}$, and $\hat{x}_{3}\in\{C_{0},C_{1}\}$, respectively. Their binary outcomes are denoted as $x_{1}\in\{\pm 1\}$, $x_{2}\in\{\pm 1\}$, and $x_{3}\in\{\pm 1\}$. We define a tripartite correlation $P(x_{1},x_{2},x_{3}|\hat{x}_{1},\hat{x}_{2},\hat{x}_{3})$ that satisfies the equality constraint 3. To illustrate genuine non-locality, consider that Alice, Bob, and Charlie share the following general pure three-qubit state,
@@
The class of pure three-qubit state that satisfies the condition (3) is of the form:
@@
We will first show that the maximum achievable value of $\mathcal{S}_{3}$ over $\mathbb{C}^{n}\otimes\mathbb{C}^{n}\otimes\mathbb{C}^{n}$ is same as over $\mathbb{C}^{2}\otimes\mathbb{C}^{2}\otimes\mathbb{C}^{2}$. Further, we will see that the quantum state $|\Psi_{max}^{\star 3}\rangle_{ABC}$ can be self-tested.
@@
$\begin{split}P(x,y,z|\hat{x},\hat{y},\hat{z})=\operatorname{Tr}[\rho(\Pi_{x|\hat{x}}\otimes\Pi_{y|\hat{y}}\otimes\Pi_{z|\hat{z}})].\end{split}$ (10)
@@
$\begin{split}\hat{x}&=(+1)\Pi_{+|\hat{x}}+(-1)\Pi_{-|\hat{x}},\quad\hat{x}\in\{A_{0},A_{1}\}\\ \hat{y}&=(+1)\Pi_{+|\hat{y}}+(-1)\Pi_{-|\hat{y}},\quad\hat{y}\in\{B_{0},B_{1}\}\\ \hat{z}&=(+1)\Pi_{+|\hat{z}}+(-1)\Pi_{-|\hat{z}},\quad\hat{z}\in\{C_{0},C_{1}\}\end{split}$ (11)